OpenStudy (anonymous):

A number divides 206 , 368 , 449 leaving the same remainder in each case. The laegest such number is?

OpenStudy (anonymous):

@Arnab09 @experimentX @apoorvk @yash2651995 @lgbasallote

OpenStudy (experimentx):

i guess it is best to start with factorizing it

OpenStudy (apoorvk):

Similar to yesterday's question. Let the number be 'n'. And the common remainder left be 'x'. Now, (206-x), (368-x), (449-x) are perfectly divisible by 'n'. Also, the difference of any of these two numbers should be divisible by 'n'. So, 449-368=81, 449-206=243, and 368-206=162, i.e. 81, 243 and 162 should be completely divisible by that 'n'. Now since, we need to find the largest possible value of 'n', a no. which divides the above there perfectly, if we find the HCF of the three nos. just above, we can find 'n'. So, HCF of 81, 162 and 243 = 9. And my friend, if you check, 9 completely satisfies every criteria in the question. So, 9 it is.